In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solution phase reaction.

A student heats 61.00 grams of lead to 99.19 °C and then drops it into a cup containing 79.31 grams of water at 24.32 °C. She measures the final temperature to be 26.53 °C.

The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.53 J/°C.

Assuming that no heat is lost to the surroundings calculate the specific heat of lead.
1
Expert's answer
2020-10-04T09:13:47-0400

Mass of lead, M = 61.00 grams

= 61× 10-3 kg

=0.061 kg

Initial temperature, T1 = 99.19 °C


After dropping it in calorimeter,

the final temperature of whole system, T2 = 26.53 °C.


Let,

heat capacity of lead = S J.kg-1.°C-1


Heat liberated by lead, H = M×S×(T2-T1)

Or, H = 0.061× S J.kg-1.°C-1 × ( 99.19 - 26.53 )°C

Or, H = 0.061×S×72.66 Joul

Or, H = 4.432×S Joul




Mass of water in the calorimeter, M' = 79.31 grams

M' = 79.31×10-3 kg

= 0.07931 kg

Initial temperature of water in calorimeter, T1' = 24.32 °C.

The final temperature of whole system, T2' = 26.53 °C.


Heat capacity of water = 1 Cal g-1 °C-1

1Cal = 4.2 Joul and 1 g = 10-3 kg

So,

Heat capacity of water, S' = 4.2J × (10-3 kg)-1 × °C-1

= 4200 J kg-1 °C-1


Heat gain by water, H' = M'×S'×(T2' - T1')

Or, H' = 0.07931 kg×4200 J kg-1 °C-1×(26.53 - 24.32) °C

Or, H' = 736.155 J

As the calorimeter is in thermal equilibrium with the water in it before and after.

So,

Initial temperature of the calorimeter = T1

= 24.32 °C


The final temperature of whole system, T2= 26.53 °C.


Heat capacity of calorimeter, S" = 1.53 J °C-1


Heat gain by the calorimeter,

H" = S"×(T2 - T1)

Or, H" = 1.53 J °C-1 × (26.53 - 24.32) °C

Or, H" = 3.381 Joul



Heat liberated by lead, H = (Heat gain by water, H' + Heat gain by the calorimeter,H")


I.e. H = H' + H"

Or, 4.432×S = 736.155 + 3.381

Or, S = (736.155 + 3.381) / 4.432

= 166.862


Hence, heat capacity of lead is 166.862 Joul kg-1 °C-1